At a temperature of 298 K, the molar enthalpy change for the conversion of red phosphorus to white phosphorus is 17.6 kJ mol-1. The molar heat capacities of red and white phosphorus are, respectively, 21.19 and 23.82 J mol-1 K-1. Calculate the molar enthalpy change at 198 K.
1
Expert's answer
2020-05-15T11:18:49-0400

According to the Kirchhoff's law, the dependence of enthalpy change of the reaction "\u2206H" on temperature "T" is related to the change in heat capacity "c_p" as follows:

"∆H_{T_2}= ∆H_{T_1} + \int_{T_1}^{T_2}{∆c_p}dT" .

The change in heat capacity for the conversion of red phosphorus to white phosphorus is:

"\u2206c_p = c_{p, white P} - c_{p, red P} = 23.82 - 21.19 =2.63" J mol-1 K-1.

Therefore, the change in enthalpy at 198 K is:

"∆H_{198 K} = ∆H_{298 K} + \int_{298}^{198}{2.63}dT"

"∆H_{198 K} = 17.6·10^{3} \text{ (J/mol)} + \int_{298}^{198}{2.63}dT"

"∆H_{198 K} = 17.6·10^{3} \text{ (J/mol)} + 2.63·(198-298)\text{ (J/mol)}"

"∆H_{198 K} = 17600 -263 = 17337 \text{ (J/mol)}" , or 17.3 kJ/mol.

Answer: the molar enthalpy change at 198 K for the conversion of red phosphorus to white phosphorus is 17.3 kJ/mol.